Calculate the surface area of a cube with sides measuring 4.7 ft.
Answer by nerdybill(7384)

Calculate the surface area of a cube with sides measuring 4.7 ft.
.
Area of each side = 4.7^2 = 22.09 square feet
.
There are 6 sides to a cube,
surface area = 6*22.09 = 132.54 square feet